OpenStudy (anonymous):

Solve. x^2 + 8x = 24 –4 ± 2 √ 2 –4 ± 2√ 10 –4 ± 2√ 5 –4 ± 2√ 7

OpenStudy (anonymous):

would it be the 2nd one?

OpenStudy (anonymous):

Move the constant term to the LHS and use the quadratic formula, or complete the square.

OpenStudy (anonymous):

I know how to do it just looking for someone to check my answer and make sure its correct

OpenStudy (anonymous):

\[x^2 + 8x = 24 \implies x^2 + 8x +16=24+16 \implies (x+4)^2=40\] \[\implies x+4= \pm \sqrt{40} \implies x=4 \pm \sqrt{40}\]

OpenStudy (anonymous):

\[4 \pm \sqrt{40} =4 \pm 2 \sqrt {10}\]

OpenStudy (anonymous):

Do math every day.

OpenStudy (anonymous):

ok so I was correct then :) thank you

OpenStudy (anonymous):

No sweat.